# Asymptotically good CSS-T codes exist

### v1

@@ 3.3 Application of transversal CCZ gate
4 Conclusion
In this paper, we introduced a construction for generating binary CSS-T codes from any given binary CSS code. This construction enabled us to prove the existence of asymptotically good binary CSS-T codes, thereby resolving an open problem proposed in [32]. Moreover, we demonstrated that our construction preserves the LDPC property of the ingredient CSS code, leading to the existence of asymptotically good binary quantum LDPC codes that are also CSS-T. Finally, we extended our construction to support transversal finer $Z$ rotations and transversal $CCZ$, and analyzed their corresponding logical operators.
@@ 4 Conclusion
Our construction, as detailed in Section 2, relies on the use of a map $\phi$ satisfying the conditions of Proposition 1. In this paper, we chose $\phi$ to be the identity map as other permutation maps induce CSS-T codes with the same parameters. An interesting remaining question is to find other alternative maps $\phi$ satisfying the conditions of Proposition 1 and leading to CSS-T codes with better parameters or new applications (see also Remark 1).
Finally, another interesting follow-up would involve the possibility of using the proposed CSS-T construction in order to design magic state distillation protocols that may improve the overhead of existing protocols. Since most of the magic state distillation protocols are related to triorthogonal codes, it would be relevant to see how more general codes such as CSS-T codes may be used for such a task. While it is noteworthy that the transversal application of some of the non-Clifford gates considered in this paper results in the logical identity, such transformations have the potential to be used for magic state distillation [25, 26]. Furthermore, it would be relevant to study how the proposed codes can yield triorthogonal codes that have a more direct relationship with current magic state distillation protocols. Since our construction yields asymptotically good codes, successfully using those for magic state distillation tasks could result in making the magic state yield parameter arbitrarily small, which is still an open question for the qubit case.
@@ Acknowledgment
This work was supported by the Spanish Ministry of Economy and Competitiveness through the MADDIE project (Grant No. PID2022-137099NBC44), by the Spanish Ministry of Science and Innovation through the project “Few-qubit quantum hardware, algorithms and codes, on photonic and solidstate systems” (PLEC2021-008251), by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call - QUANTUM SPAIN project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda, and by the grant ANR-21-CE39-0009-BARRACUDA from the French National Research Agency. J.E.M. is funded by the Spanish Ministry of Science, Innovation and Universities through a Jose Castillejo mobility grant for his stay at the Cavendish Laboratory of the University of Cambridge.

### v2

@@ IV-C Application of transversal CCZ gate
V New construction of triorthogonal codes
In this section, we discuss a method for transforming certain CSS-T codes into triorthogonal codes. Then, we apply this result to provide a general construction of triorthogonal codes using binary self-orthogonal codes (Theorem 12), which is the main result of this section. Finally, we provide examples of triorthogonal codes constructed using our approach.
Since triorthogonal codes are a subclass of CSS-T codes, it is natural to search for a method of transforming CSS-T codes into them. This would allow using the magic state distillation protocol discussed in [11] for the transformed codes and find new applications of CSS-T codes. The following theorem gives an instance of such transformations.
Theorem 11.
Let $C_{2}\subseteq C_{1}\subseteq\mathbb{F}_{2}^{n}$ and $C_{1}=C_{2}\oplus{\textrm{S}pan}\{w_{1},\ldots,w_{m}\}$ where each $w_{i}$ has an odd weight. Suppose that ${\textrm{S}pan}\{w_{1},\ldots,w_{r}\}$ forms a triorthogonal subspace for some $r\leq m$. If $(C_{1},C_{2})$ is a CSS-T pair , then $C_{2}^{\prime}=C_{2}\subseteq C_{1}^{\prime}=C_{2}\oplus{\textrm{S}pan}\{w_{1},\ldots,w_{r}\}$ forms a triorthogonal CSS code.
Proof.
Since $(C_{1},C_{2})$ is a CSS-T pair, we have $C_{2}^{\star 2}\subseteq C_{2}^{\bot}$, implying that $C_{2}$ is a triorthogonal subspace. For $j\leq r$, we have the following sequence of implications for each $c_{2}$ and $c_{2}^{\prime}\in C_{2}$:
$\begin{split}&C_{1}^{\star 2}\subseteq C_{2}^{\perp}\implies(w_{j}\star c_{2})\cdot c_{2}^{\prime}=0\pmod{2}\\ &\implies{\textrm{w}t}(w_{j}\star c_{2}\star c_{2}^{\prime})=0\pmod{2}.\end{split}$
Taking $c_{2}=c_{2}^{\prime}$, we get $\forall c_{2}\in C_{2},{\textrm{w}t}(w_{j}\star c_{2})=0\pmod{2}$. On the other hand, the following holds as well for each $c_{2}\in C_{2}$ and $j,k\leq m$:
$C_{1}^{\star 2}\subseteq C_{2}^{\perp}\implies(w_{j}\star w_{k})\cdot c_{2}=0\implies{\textrm{w}t}(w_{j}\star w_{k}\star c_{2})=0\pmod{2}.$
Using the linearity property of these codes, one can extend the above results to any arbitrary vectors in $C_{1}$ and $C_{2}$, which proves the triorthogonality of the CSS code constructed from $C_{2}\subseteq C_{1}$. ∎
Corollary 4.
Let $C_{2}\subseteq C_{1}\subseteq\mathbb{F}_{2}^{n}$ form a CSS-T pair. Then the CSS code of $C_{1}^{\prime}=C_{2}\oplus{\textrm{S}pan}\{w\}$, where $w\in C_{1}$ is a single vector of odd weight, and $C_{2}^{\prime}=C_{2}$ is a triorthogonal code.
Proof.
The result follows immediately since a single vector forms a triorthogonal subspace. ∎
Through the last part of this section, we generalize the doubling construction of triorthogonal codes to expand the spectrum of suitable codes for magic state distillation. The doubling transformation was first discussed in [38] and has been used to generate triply-even CSS codes, a class of codes closely related to triorthogonal codes but with some subtle differences. Recently, several triply-even CSS codes with optimal parameters have been constructed using this approach [2]. The main ingredients in the doubling construction are a doubly-even code and a triorthogonal code. In the following theorem, we relax the previous requirement by using a pair of self-orthogonal and triorthogonal codes. This generalizes the earlier construction by enabling the production of triorthogonal codes that are not necessarily triply-even. We represent the all-one and all-zero vectors of length $n$ by $\bf{1_{n}}$ and $\bf{0_{n}}$, respectively.
Theorem 12.
Let $C_{2}$ be a binary self-orthogonal code of an odd length $n_{1}$ and $C_{1}=C_{2}\oplus{\textrm{S}pan}\{\bf{1_{n_{1}}}\}$ with the corresponding CSS parameters $[[n_{1},1,d_{1}]]$. Let $C^{\prime}_{2}\subset C_{1}^{\prime}$ be a triorthogonal CSS code with parameters $[[n_{2},1,d_{2}]]$, where $n_{2}$ is also an odd integer. Then one can construct a triorthogonal code, not necessarily triply-even, with parameters $[[2n_{1}+n_{2},1,\min\{d_{1},d_{2}+2\}]]$.
Proof.
Without loss of generality, we assume that $C_{1}^{\prime}=C_{2}^{\prime}\oplus{\textrm{S}pan}\{\bf{1_{n_{2}}}\}$ as otherwise the same proof works by replacing $\bf{1_{n_{2}}}$ with other relevant vector. One can construct a pair of new binary codes, namely $C_{2}^{\prime\prime}\subset C_{1}^{\prime\prime}$, where $C_{2}^{\prime\prime}$ has the generator matrix
$G=\begin{bmatrix}C_{2}&C_{2}&\bf{0_{n_{2}}}\\ \bf{0_{n_{1}}}&\bf{0_{n_{1}}}&C_{2}^{\prime}\\ \bf{0_{n_{1}}}&\bf{1_{n_{1}}}&\bf{1_{n_{2}}}\end{bmatrix},$
and $C_{1}^{\prime\prime}=C_{2}^{\prime\prime}\oplus{\textrm{S}pan}\{\bf{1_{2n_{1}+n_{2}}}\}$. The CSS code of $C_{2}^{\prime\prime}\subseteq C_{1}^{\prime\prime}$ is a $[[2n_{1}+n_{2},1,d]]$ code. Moreover, a similar proof as that of [39, Lemma 3] can be employed to show that $d=\min\{d_{1},d_{2}+2\}$. Next, we show that $C_{2}^{\prime\prime}\subseteq C_{1}^{\prime\prime}$ is a CSS-T pair. Since $C_{2}$ is a self-orthogonal code of an odd length, we have ${\bf{1_{n_{1}}}}\in C_{2}^{\bot}$ and $x\star y$ has an even weight for each $x,y\in C_{2}$. Then for each $u$ and $v\in C_{2}^{\prime\prime}$, the vector $u\star v$ has one of the forms
  1. 1.
    $u\star v=(x\star y,x\star y,{\bf{0_{n_{2}}}})$ for some $x,y\in C_{2}$,
  2. 2.
    $u\star v=({\bf{0_{n_{1}}}},{\bf{0_{n_{1}}}},x^{\prime}\star y^{\prime})$ for some $x,y\in C_{2}^{\prime}$,
  3. 3.
    $u\star v=({\bf{0_{n_{1}}}},{\bf{1_{n_{1}}}},{\bf{1_{n_{2}}}})$.
  4. 4.
    $u\star v=({\bf{0_{n_{1}}}},{x},{\bf{0_{n_{2}}}})$ for some $x\in C_{2}$,
or a linear combination of them. Moreover, an arbitrary codeword of $C_{2}^{\prime\prime}$ is in the form of
$w=(a_{1}x_{1},a_{1}x_{1}+a_{3}{\bf{1_{n_{1}}}},a_{2}x_{1}^{\prime}+a_{3}{\bf{1_{n_{2}}}}),$
where $x_{1}\in C_{2}$, $x_{1}^{\prime}\in C_{2}^{\prime}$, and $a_{i}\in\mathbb{F}_{2}$ for $i=1,2,3$. An easy observation based on the forms (1)–(4) and the structure of $w$ shows that $(u\star v)\cdot w=0\pmod{2}$ for any $u,v,w\in C_{2}^{\prime\prime}$. Hence, we have $C_{2}^{\prime\prime}\star C_{2}^{\prime\prime}\subseteq C_{2}^{\prime\prime\bot}$. Moreover, ${\bf{1_{2n_{1}+n_{2}}}}\star{\bf{1_{2n_{1}+n_{2}}}}={\bf{1_{2n_{1}+n_{2}}}}\in C_{2}^{\prime\prime\bot}$, and for any $w\in C_{2}^{\prime\prime}$, we have $w\star{\bf{1_{2n_{1}+n_{2}}}}=w\in C_{2}^{\prime\prime\bot}$. Hence we conclude that $C_{1}^{\prime\prime}\star C_{1}^{\prime\prime}\subseteq C_{2}^{\prime\prime\bot}$ or equivalently $C_{2}^{\prime\prime}\subseteq C_{1}^{\prime\prime}$ is a CSS-T pair. Finally, such CSS-T pair satisfies the conditions of Corollary 4, hence it is a triorthogonal code. ∎
Theorem 12 allows for the construction of triorthogonal codes more easily, and enables a broader spectrum of classical codes to be used for this purpose. All the triorthogonal codes discussed in [2] can be obtained using the above theorem. Moreover, one can get codes beyond them. In the next section, we discuss how to extract suitable ingredients from the family of binary self-dual codes, following the strategy used in [2], but focusing exclusively on singly even codes.
Using this approach, one gets a broader list of codes for constructing triorthogonal codes as: a binary doubly-even self-dual code of length $n$ satisfies $n\equiv 0\pmod{8}$ [40, Page 31]. However, a binary singly even self-dual code of length $n$ exists for each even integer $n$. Moreover, the number of singly even codes is much larger than the doubly-even codes. For instance, the number of singly-even (respectively, doubly-even) codes of lengths 24 and 32 are 46 (respectively, 9) and 3210 (respectively, 85), respectively.
V-A Ingredient codes and examples
In practice, obtaining a triorthogonal code appears to be more challenging than obtaining a binary self-orthogonal code. Therefore, to apply the result of Theorem 12, it is more efficient to fix a triorthogonal code and search for quantum self-orthogonal codes. So, in what follows, we discuss a scenario for constructing quantum self-orthogonal codes.
Note that all the computations displayed in the following examples, including the actual minimum distances of our codes, were done using the Magma computer algebra system [41].
Let $C$ be a binary self-dual code of length $n$ (which is thus an even integer) with minimum distance at least 2. Then $C=C^{\bot}$ and has parameters $[n,\frac{n}{2}]$. Shortening the code $C$ in any position gives $C_{2}$, which is a self-orthogonal code and it has parameters $[n-1,\frac{n}{2}-1]$. Then choosing $C_{1}=C_{2}\oplus{\textrm{S}pan}\{\bf{1_{n-1}}\}$ leads to a pair $C_{2}\subseteq C_{1}$ which is an $[[n-1,1]]$ CSS code. One can easily see that such a code satisfies the conditions of Theorem 12.
This approach replaces the requirement for a binary self-orthogonal code in Theorem 12 with a new ingredient, namely a binary self-dual code. Self-dual codes form an important class of classical codes, and several best-known codes and families of quantum codes have been constructed using them. For instance, the parameters of several self-dual codes are calculated in [42] (until length 130) and in [43] (until length 241 mix of binary and quaternary). Moreover, there exist other databases of self-dual codes with good parameters, for instance [44, 45], that allow easier integration of such codes to build triorthogonal codes. Next, we produce some examples of triorthogonal codes using the approach discussed above.
Applying the result of Theorem 12 to the triorthogonal code $[[15,1,3]]$ (see [39]), and the singly even self-dual codes of lengths $18$ and $20$ in [45] gives triorthogonal codes with parameters
  1. 1.
    $[[49,1,3]]$ (degenerate to 2) and $[[49,1,5]]$ (degenerate to 4 and it was discussed in [39]),
  2. 2.
    $[[53,1,3]]$ (non-degenerate and degenerate to 2) and $[[53,1,5]]$ (degenerate to 4).
In the following example, we consider a case where a singly even code yields a better triorthogonal code than those obtained from doubly-even codes with the same minimum distance.
Example 1.
Let $C$ be the linear binary cyclic code of length $89$ generated by the polynomial
$\begin{split}g(x)=&x^{45}+x^{44}+x^{42}+x^{38}+x^{36}+x^{35}+x^{33}+x^{32}+\\ &x^{30}+x^{27}+x^{26}+x^{24}+x^{23}+x^{20}+x^{19}+x^{18}+\\ &x^{16}+x^{15}+x^{12}+x^{8}+x^{5}+x^{4}+x^{3}+1.\end{split}$
One can verify that $C$ is self-orthogonal and $C^{\bot}=C\oplus{\textrm{S}pan}\{{\bf{1_{89}}}\}$. Moreover, $C$ is singly even and cannot be used as an ingredient in the triorthogonal construction of [2]. Our computation shows that the CSS code of $C\subseteq C^{\bot}$ has parameters $[[89,1,17]]$ (degenerate to 12). Then applying Theorem 12 to this code and the distance 15 triorthogonal code of [2], namely $[[575,1,15]]$, produces a triorthogonal code with parameters $[[753,1,17]]$ which has better parameters than the minimum distance 17 triorthogonal code produced in [2], namely $[[777,1,17]]$.
VI Conclusion
In this paper, we proved the existence of asymptotically good CSS-T quantum (LDPC) codes, resolving an open problem proposed in [16]. We also showed that the same holds for other finer $Z$ rotations. Our main technique was based on employing a construction of CSS-T codes integrated from concatenation and a mapping $\phi$. In the particular case of $\phi=I$, our technique overlapped that of [1]. For the codes obtained from our approach, we also analyzed the logical operator of certain transversal gates. Particularly, we concluded that such logical operator is always the identity, and discussed the application of such codes for mitigating coherent noise.
We also introduced a novel doubling transformation that extends the doubling construction described in [2] for generating triorthogonal codes. Our approach accommodates self-orthogonal codes as the main component, instead of the doubly-even codes discussed in [2, 39]. This generalization expands the class of codes suitable for constructing triorthogonal codes, thereby offering greater flexibility for magic state distillation.
@@ VI Conclusion
In this paper, we chose the map $\phi$ to be the identity map as other permutation maps induce CSS-T codes with the same parameters. An interesting remaining question is to find other alternative maps $\phi$ satisfying the conditions of Proposition 1 and leading to CSS-T codes with better parameters or new applications (see also Remark 1). As another future direction, we aim to systematically search for non-doubly-even self-orthogonal codes to serve as the main ingredient in our triorthogonal construction. This can be achieved through a study of certain classical codes and their parameters, with the goal of potentially reducing the overhead of magic state distillation using such codes.
@@ Acknowledgment
The order of the authors is purely alphabetical. All authors contributed to the paper to the same extent.
The authors thank Shubham P. Jain for an insightful private communication. They are also grateful to Narayanan Rengaswamy for his support and valuable advice that led to the final version of the paper.